A Man of mass 80 kg dives into a swimming pool from a tower of height 18 m. He was found to go down in water by 2.2 m and then started rising. Find the average resistance of water. Neglect the air resistance of water take value of, g =9.81m/s2
1
Expert's answer
2019-12-20T04:48:18-0500

Considering the jump to the pool,


"v^2 = u^2 + 2gh \newline
where \space u=0 \newline
h= height \space of \space the \space tower \newline
v= final \space velocity"



"g =9.81m s^{-2} \newline
\because v=\sqrt{2gh} \newline
v=\sqrt{2*9.81*18} \newline
=\underline{\underline{18.79ms^{-1} }} \\~\\

This \space v \space acts \space as \space u' \space for \space deaccelaration \space in \space the \space water \\~\\

For \space the \space water, \newline
v'^2 = u'^2 -2as \newline
where\space v'=0 , \newline
u'=18.79ms^{-1} \newline
s=2.2m \\~\\
solving \space this \space gives \\~\\
a=\underline{\underline{80.24ms^{-2}}} \\~\\ Substituding \space to \space F=ma \newline
where \space m= 80kg \newline
F=80*80.24 \newline
=\underline{\underline{6419.2N}}"